# Prediction of turbulent channel flow using Fourier neural operator-based machine-learning strategy

### v1

@@ Abstract
Fast and accurate predictions of turbulent flows are of great importance in the science and engineering field. In this paper, we investigate the implicit U-Net enhanced Fourier neural operator (IUFNO) in the stable prediction of long-time dynamics of three-dimensional (3D) turbulent channel flows. The trained IUFNO models are tested in the large-eddy simulations (LES) at coarse grids for three friction Reynolds numbers: $Re_{\tau}\approx 180$, $395$ and $590$. The adopted near-wall mesh grids are tangibly coarser than the general requirements for wall-resolved LES. The numerical experiments show that the IUFNO framework outperforms the traditional dynamic Smagorinsky model (DSM) and the wall-adapted local eddy-viscosity (WALE) model in the predictions of a variety of flow statistics and structures, including the mean and fluctuating velocities, the probability density functions (PDFs) and joint PDF of velocity fluctuations, the Reynolds stress profile, the kinetic energy spectrum, and the Q-criterion (vortex structures). Meanwhile, the trained IUFNO models are computationally much faster than the traditional LES models. Thus, the IUFNO is a promising approach for the fast prediction of wall-bounded turbulent flow.
@@ I Introduction
In the recent decade, machine-learning-based flow prediction methods have been extensively proposed due to the fast development of modern computers and the accumulation of high fidelity data[22, 24, 25, 23]. These attempts include, but not limited to, learning the important aerodynamic forces in the flow field such as the lifting and dragging force coefficient [26, 27, 28], reconstructing part of governing equations such as the modeling of the RANS and LES closure terms [30, 31, 29, 32, 33, 34, 40, 39, 41, 43, 36, 37, 38, 35, 42], wall modeling [44, 45], inferring the missing information in the case of measurement constraints or flow field damage [47, 48, 46, 49], flow field super-resolution [50, 51], and directly predicting the evolution of the flow field [54, 52, 53, 55, 56, 59, 60, 61, 58, 62, 63, 57]. Among these applications, directly predicting the temporal evolution of the flow field has attracted increasing attentions these days, since the trained model does not require to solve the Navier-stokes equations while giving a fast evaluation of the detailed flow information. The adopted methods mainly include the recursive neural network (RNN) and long short term memory (LSTM)-based frameworks [59, 60, 64, 65, 66, 61, 58], physics-informed neural networks (PINN)-based methods [67, 68, 69], and neural operator-based methods [54, 52, 70, 55, 56]. For example, Bukka et al. combined the proper orthogonal decomposition (POD) and deep learning approach to predict the flow past a cylinder [58]. Han et al. conducted a series of studies using the CNN and LSTM-based framework in the predictions of flow field and fluid-solid interactions [64, 65, 66]. Raissi et al. developed a PINN-based framework to predict the solutions of general nonlinear partial differential equations [67]. In the two-dimensional (2D) Rayleigh Bénard convection problem, Wang et al. developed a turbulent flow network (TF-Net) based on the specialized U-Net with incorporated physical constraints [68].
@@
While many neural networks (NN) are effective at approximating the mappings between finite-dimensional Euclidean spaces for a given data set, these NN-based models are difficult to generalize to different flow conditions or boundary conditions [67, 52]. In this consideration, Li et al. proposed a novel Fourier neural operator (FNO) framework which can efficiently learn the mappings between the high dimensional features in Fourier space, and it enables reconstructing the information in infinite dimensional spaces [52]. In their numerical tests, the proposed FNO model achieves outstanding accuracy in the prediction of two-dimensional (2D) turbulence. Since Li et al., many extensions and applications of FNO have emerged [70, 53, 75, 76, 55, 56, 77, 54, 78, 71, 72, 73, 74, 79]. Lehmann et al. applied the FNO in the prediction of the propagation of seismic waves [76]. To improve the prediction accuracy of FNO for 2D turbulence at high Reynolds numbers, Peng et al. introduced the attention mechanism into the FNO framework, resulting in better statistical properties and instantaneous structures of the flow field [55]. Wen et al. combined FNO with U-Net to solve the complex gas-liquid multiphase problems with improved accuracy compared to the original FNO and the CNN frameworks [77]. You et al. proposed an implicit Fourier neural operator (IFNO) [79], which can greatly reduce the number of trainable parameters and memory cost compared to multi-layer structure of the FNO [75].
@@ IV.1 The DNS database and the configuration of the training data set
The training setups are configured as follows: For both the FNO and IUFNO models, two fully-connected layers (i.e. $P$ and $Q$ in Fig. 1) are configured before and after the Fourier layer(s). For the FNO model, four consecutive Fourier layers are adopted [70]. For the IUFNO, the consecutive Fourier layers are replaced by an implicit Fourier layer with embedded U-Net structure, and the numbers of internal iterations for the Fourier layer in the IUFNO are 40, 20 and 20 for $Re_{\tau}\approx 180$, $395$ and $590$, respectively. In the U-Net structure of the IUFNO model, three consecutive encoders are configured, followed by three decoders with skip connections [86]. For all the trainings, the cutoff wavenumbers for the Fourier modes truncation are based on the half of the minimum grid numbers (i.e. in z direction). Hence, we set $k_{max}=8$, $16$ and $16$ for $Re_{\tau}\approx 180$, $395$ and $590$, respectively. The Adam optimizer is used for optimization [92], the initial learning rate is set to $10^{-3}$, and the GELU function is chosen as the activation function [93]. The training and testing losses are defined as
[…]
Here, $\Delta u^{*}$ denotes the prediction of velocity increments and $\Delta u$ is the ground truth. The training and testing curves at $Re_{\tau}\approx 180$ for the FNO and IUFNO models are shown in Figs. 3a and 3b, respectively. As can be seen, the IUFNO has a lower loss compared to the FNO model.
Figure 3: The evolutions of the loss curves: (a) FNO; (b) IUFNO.
@@
In the a posteriori tests, the initial turbulent fields are taken from a new fDNS field which is different from those in the training set, such that we can test whether the trained model can be generalize to different initial conditions. Meanwhile, the no-slip condition is reinforced in FNO and IUFNO in the a posteriori tests. To compare the performance of the FNO and IUFNO frameworks, we show the predicted mean streamwise velocity profiles at different time instants for $Re_{\tau}\approx 180$ in Fig. 4. It can be seen that both the FNO and the IUFNO predict well the velocity profile initially. However, as the solutions evolve, the FNO result diverges from the fDNS benchmark, while the IUFNO remains adequately accurate. At $t=600\Delta T$ and $800\Delta T$, both of which are far beyond the time range of the training set ($t<400\Delta T$), the IUFNO remains stable and accurate. In physical sense, $800\Delta T$ allows the bulk flow to pass through the channel for approximately 42.5 times, demonstrating the long-term predictive ability of the IUFNO model. Hence, in the rest of the work, only the results of IUFNO will be presented and compared against the traditional LES models including the DSM and WALE models.
Figure 4: The mean streamwise velocities at $Re_{\tau}\approx 180$: (a) $t=10\Delta T$; (b) $t=15\Delta T$; (c) $t=20\Delta T$; (d) $t=25\Delta T$; (e) $t=600\Delta T$; (f) $t=800\Delta T$.
The training and testing losses for $Re_{\tau}\approx 395$ and $590$ are shown in Figs. 5a and 5b, respectively. It can be seen that as $Re_{\tau}$ increases, the loss goes up, indicating the increased difficulty in learning high Reynolds number turbulent flow. Meanwhile, to avoid over-fitting of the model, we have applied early stopping which is widely used in gradient descent learning [94, 95]. In both cases, the model parameters are extracted after 30 epochs before the performance on the testing set starts to degrade.
@@
In Figs. 6, 7 and 8, the mean streamwise velocity and root-mean-squared (rms) fluctuating velocities predicted by the IUFNO model are displayed for $Re_{\tau}\approx 180$, $395$ and $590$, respectively. Also shown in the figures are the corresponding predictions by the DSM and WALE models. As can be seen, the IUFNO model performs reasonably better at all three Reynolds numbers compared to the DSM and WALE models. Nevertheless, the prediction of the mean velocity by IUFNO at $Re_{\tau}\approx 590$ is not as good as that at $Re_{\tau}\approx 180$ and $395$. This also reflects the increasing difficulty in predicting turbulent flows at high Reynolds number. For the traditional LES models, we observe that the WALE model performs much better than the DSM. Meanwhile, the WALE model diverges at $Re_{\tau}\approx 590$, thus the results of WALE are not presented for $Re_{\tau}\approx 590$. Here, it should be emphasized that the current LES grids are coarser than the general requirements for wall-resolved LES [5, 91]. In the expense of additional computational cost (i.e. using sufficiently fine grids or carefully configured wall model), the traditional LES models can also achieve reasonably accurate LES predictions, which is not the concern of the current work.
@@
The probability density functions (PDFs) of the three fluctuating velocity components are shown in Fig. 9 for different friction Reynolds numbers. The PDFs of the velocity fluctuations are all symmetric, and the ranges of the PDFs are in consistency with the corresponding rms values in Figs. 6, 7 and 8. It can be seen that the IUFNO model gives reasonably good predictions of the PDFs at all three Reynolds numbers. Notably, the advantage of the IUFNO model is more pronounced for the streamwise components of the fluctuating velocity compared to the normal and spanwise components whose predictions by the WALE model are also satisfying. Again, the DSM gives the worst predictions among all models.
Figure 9: The probability density functions (PDFs) of the velocity components at different Reynolds numbers: (a) $u^{\prime},Re_{\tau}\approx 180$; (b) $v^{\prime},Re_{\tau}\approx 180$; (c) $w^{\prime},Re_{\tau}\approx 180$; (d) $u^{\prime},Re_{\tau}\approx 395$; (e) $v^{\prime},Re_{\tau}\approx 395$; (f) $w^{\prime},Re_{\tau}\approx 395$; (g) $u^{\prime},Re_{\tau}\approx 590$; (h) $v^{\prime},Re_{\tau}\approx 590$; (i) $w^{\prime},Re_{\tau}\approx 590$.
The predicted shear Reynolds stresses by different LES models are shown in Fig. 10. The maximum shear Reynolds stresses are located near the upper and lower walls where both the mean shear effects and the velocity fluctuations are strong. The distribution between the two peaks is approximately linear, which is consistent with the literature [96]. Both the IUFNO and WALE models can predict the Reynolds stress well at $Re_{\tau}\approx 180$ and $395$ while significant discrepancy can be observed for the DSM. At $Re_{\tau}\approx 590$, the IUFNO model still performs better than the DSM, but it has some wiggles in the linear region.
@@
As shown in Fig. 11, at all three Reynolds numbers, the joint PDFs close to the wall are much wider in the second and fourth quadrants, indicating that shear events (ejection and sweep) are more dominant near the wall rather than the center of the channel. Meanwhile, we also observe that the IUFNO model has very close agreements with the fDNS results. The predictions by the WALE model are similar to those by the IUFNO model, while the predictions by the DSM are relatively worse.
@@
To examine the energy distribution at different scales, we calculate the kinetic energy spectrum in the LES. The streamwise spectra are shown in Figs. 12a, 12b and 12c for $Re_{\tau}\approx 180$, $395$ and $590$, respectively. At $Re_{\tau}\approx 180$, the predicted spectrum by the IUFNO model agrees well with the fDNS result. Both the DSM and WALE models overestimate the energy spectrum while the WALE model outperforms the DSM. This is also consistent with the results for the rms fluctuating velocity magnitude in Fig. 6. At $Re_{\tau}\approx 395$, the spectrum predicted by IUFNO slightly deviates from the fDNS result, but it still outperforms the DSM and WALE models. For the spectrum at $Re_{\tau}\approx 590$, the IUFNO model still outperforms the DSM, however, some nonphysical discontinuities in the IUFNO spectrum can be observed. The cause of such phenomenon is yet unclear.
The spanwise kinetic energy spectra are shown in Figs. 12d, 12e and 12f for $Re_{\tau}\approx 180$, $395$ and $590$, respectively. Unlike the streamwise spectrum, the maximum energy in the spanwise spectrum is not contained in the first wavenumber, presumably because there are no dominant largest-scale motions in the spanwise direction. Overall, the IUFNO has a closer agreement with the fDNS result compared to the other models.
@@
where $\overline{\Omega}_{ij}=\frac{1}{2}(\partial{\overline{u}_{i}}/\partial{x_{j}}-\partial{\overline{u}_{j}}/\partial{x_{i}})$ is the filtered rotation rate. The instantaneous iso-surfaces of $Q$ for $Re_{\tau}\approx 180$, $395$ and $590$ are displayed in Figs. 13, 14 and 15, respectively. The iso-surfaces are colored by the streamwise velocities. As the figures depict, the vortex structures predicted by the DSM are denser than the fDNS structures, and the streamwise velocities are larger than the corresponding fDNS values. In contrast, the WALE and IUFNO models yield better predictions of the vortex structures. Hence, the ability of the IUFNO model to predict the vortex structures is confirmed.
@@
Finally, we compare the computation cost in the LES for different models in Table 3. The values in the table are the time consumption (in seconds) for equivalent 10000 DNS time steps. The IUFNO models are tested on the NVIDIA A100 GPU, where the CPU type is AMD EPYC 7763 @2.45GHz. The DSM and WALE simulations are performed on the CPU, which is Intel Xeon Gold 6148 @2.40 GHz. Considering the number of computation cores (shown in the parentheses) used by the DSM and WALE models, the computational efficiency of the IUFNO framework is considerably higher than those of the traditional LES models.
$Re_{\tau}$ Cost of DSM Cost of WALE Cost of IUFNO
180 97.4s ($\times$16 cores) 52.1s ($\times$16 cores) 12.5s
395 287.2s ($\times$32 cores) 146.3s ($\times$32 cores) 18.1s
590 315.3s ($\times$64 cores) NA 19.6s

### v4

@@ Abstract
Fast and accurate predictions of turbulent flows are of great importance in the science and engineering field. In this paper, we investigate the implicit U-Net enhanced Fourier neural operator (IUFNO) in the stable prediction of long-time dynamics of three-dimensional (3D) turbulent channel flows. The trained IUFNO models are tested in the large-eddy simulations (LES) at coarse grids for three friction Reynolds numbers: $Re_{\tau}\approx 180$, $395$ and $590$. The adopted near-wall mesh grids are tangibly coarser than the general requirements for wall-resolved LES. Compared to the original Fourier neural operator (FNO), the implicit FNO (IFNO) and U-Net enhanced FNO (UFNO), the IUFNO model has a much better long-term predictive ability. The numerical experiments show that the IUFNO framework outperforms the traditional dynamic Smagorinsky model (DSM) and the wall-adapted local eddy-viscosity (WALE) model in the predictions of a variety of flow statistics and structures, including the mean and fluctuating velocities, the probability density functions (PDFs) and joint PDF of velocity fluctuations, the Reynolds stress profile, the kinetic energy spectrum, and the Q-criterion (vortex structures). Meanwhile, the trained IUFNO models are computationally much faster than the traditional LES models. Thus, the IUFNO model is a promising approach for the fast prediction of wall-bounded turbulent flow.
@@ I Introduction
In the recent decade, machine-learning-based flow prediction methods have been extensively proposed due to the fast development of modern computers and the accumulation of high fidelity data[22, 24, 25, 23, 26, 27]. These attempts include, but not limited to, learning the important aerodynamic forces in the flow field such as the lifting and dragging force coefficient [28, 29, 30], reconstructing part of governing equations such as the modeling of the RANS and LES closure terms [32, 33, 31, 34, 35, 36, 42, 41, 43, 45, 38, 39, 40, 37, 44], wall modeling [46, 47], inferring the missing information in the case of measurement constraints or flow field damage [49, 50, 48, 51], flow field super-resolution [52, 53], and directly predicting the evolution of the flow field [56, 54, 55, 57, 58, 61, 62, 63, 60, 64, 65, 59, 66, 67, 68, 69]. Among these applications, directly predicting the temporal evolution of the flow field has attracted increasing attentions these days, since the trained model does not require to solve the Navier-stokes equations while giving a fast evaluation of the detailed flow information. The adopted methods mainly include the recursive neural network (RNN) and long short term memory (LSTM)-based frameworks [61, 62, 71, 72, 73, 63, 60], physics-informed neural networks (PINN)-based methods [74, 75, 76], and neural operator-based methods [56, 54, 77, 57, 58]. For example, Bukka et al. combined the proper orthogonal decomposition (POD) and deep learning approach to predict the flow past a cylinder [60]. Han et al. conducted a series of studies using the CNN and LSTM-based framework in the predictions of flow field and fluid-solid interactions [71, 72, 73]. Raissi et al. developed a PINN-based framework to predict the solutions of general nonlinear partial differential equations [74]. In the two-dimensional (2D) Rayleigh Bénard convection problem, Wang et al. developed a turbulent flow network (TF-Net) based on the specialized U-Net with incorporated physical constraints [75].
@@
While many neural networks (NN) are effective at approximating the mappings between finite-dimensional Euclidean spaces for a given data set, these NN-based models are difficult to generalize to different flow conditions or boundary conditions [74, 54]. In this consideration, Li et al. proposed a novel Fourier neural operator (FNO) framework which can efficiently learn the mappings between the high dimensional features in Fourier space, and it enables reconstructing the information in infinite dimensional spaces [54]. In their numerical tests, the proposed FNO model achieves outstanding accuracy in the prediction of two-dimensional (2D) turbulence. Since Li et al., many extensions and applications of FNO have emerged [77, 55, 82, 83, 57, 58, 84, 56, 85, 78, 79, 80, 81, 86, 87, 90, 88, 89]. Lehmann et al. applied the FNO in the prediction of the propagation of seismic waves [83]. To improve the prediction accuracy of FNO for 2D turbulence at high Reynolds numbers, Peng et al. introduced the attention mechanism into the FNO framework, resulting in better statistical properties and instantaneous structures of the flow field [57]. Wen et al. developed a U-Net enhanced FNO (UFNO) to solve the complex gas-liquid multiphase problems with improved accuracy compared to the original FNO and the CNN frameworks [84]. You et al. proposed an implicit Fourier neural operator (IFNO) [86], which can greatly reduce the number of trainable parameters and memory cost compared to multi-layer structure of the FNO [82].
@@ IV.1 The DNS database and the configuration of the training data set
The training setups are configured as follows: For both the FNO and IUFNO models, two fully-connected layers (i.e. $P$ and $Q$ in Fig. 1) are configured before and after the Fourier layer(s). For the FNO model, four consecutive Fourier layers are adopted [77]. For the IUFNO model, the consecutive Fourier layers are replaced by an implicit Fourier layer with embedded U-Net structure, and the numbers of internal iterations for the Fourier layer in the IUFNO model are 40, 20 and 20 for $Re_{\tau}\approx 180$, $395$ and $590$, respectively. In the U-Net structure of the IUFNO model, three consecutive encoders are configured, followed by three decoders with skip connections [97]. For all the trainings, the cutoff wavenumbers for the Fourier modes truncation are based on the half of the minimum grid numbers (i.e. in z direction). Hence, we set $k_{max}=8$, $16$ and $16$ for $Re_{\tau}\approx 180$, $395$ and $590$, respectively. The Adam optimizer is used for optimization [103], the initial learning rate is set to $10^{-3}$, and the GELU function is chosen as the activation function [104]. The training and testing losses are defined as
$\textit{Loss}=\frac{\|\Delta u^{*}-\Delta u\|_{2}}{\|\Delta u\|_{2}},\text{ where }\|\mathbf{A}\|_{2}=\frac{1}{n}\sqrt{\sum_{k=1}^{n}\left|\mathbf{A}_{\mathbf{k}}\right|^{2}}.$ (18)
Here, $\Delta u^{*}$ denotes the prediction of velocity increments and $\Delta u$ is the ground truth. The evolution curves of the training and testing losses at $Re_{\tau}\approx 180$ are shown in Fig. 3. Here, we also present the corresponding results for the implicit FNO (IFNO) [86] which does not incorporate the U-Net structure, and the results for the U-Net enhanced FNO (UFNO) [84] which uses consecutive U-Net embedded Fourier layers instead of the implicit looping structure. As can be seen, the IUFNO model has the lowest loss compared to the other FNO models. We also observe that the U-Net-based models (UFNO and IUFNO) converge faster compared to the other models.
Figure 3: The evolutions of the loss curves: (a) FNO; (b) IFNO; (c) UFNO; (d) IUFNO.
@@
In the a posteriori tests, the initial turbulent fields are taken from a new fDNS field which is different from those in the training set, such that we can test whether the trained model can be generalize to different initial conditions. Meanwhile, the no-slip condition is reinforced in FNO and IUFNO in the a posteriori tests. To compare the performances of different FNO-based frameworks, we show the predicted mean streamwise velocity profiles at different time instants for $Re_{\tau}\approx 180$ in Fig. 4. It can be seen that all the FNO models predict the velocity profile well initially (at $t=10\Delta T$). However, as the solutions evolve, the predictions of the FNO, IFNO and UFNO models diverge from the fDNS benchmark sooner or later, while the IUFNO model remains adequately accurate. At $t=25\Delta T$, the FNO result starts to strongly deviate from the fDNS benchmark, while the other models are still reasonably accurate. At $t=29\Delta T$, FNO diverges, and UFNO strongly deviates from the fDNS result, while IFNO only slightly deviates. At $t=40\Delta T$, all models diverge except the IUFNO model. Most importantly, at $t=600\Delta T$ and $800\Delta T$, both of which are far beyond the time range of the training set ($t<400\Delta T$), the IUFNO model remains numerically stable and accurate. In physical sense, $800\Delta T$ allows the bulk flow to pass through the channel for approximately 42.5 times, demonstrating the long-term predictive ability of the IUFNO model. Hence, in the rest of the work, only the results of IUFNO will be presented and compared against the traditional LES models including the DSM and WALE models.
Figure 4: The mean streamwise velocities at $Re_{\tau}\approx 180$ for different FNO-based LES models: (a) $t=10\Delta T$; (b) $t=25\Delta T$; (c) $t=29\Delta T$; (d) $t=40\Delta T$; (e) $t=600\Delta T$; (f) $t=800\Delta T$.
Since the IUFNO framework is trained on DNS data, it is interesting to study how the model converges with different amount of training data. Fig. 5 displays the evolutions of the testing loss with different amount of training data at $Re_{\tau}=180$. As can be seen, when the amount of training data is small (5 and 10 groups of DNS), the testing losses only decrease to the level of 0.2 to 0.25. As the training data set increases to 15 groups, a significant reduction of the testing loss can be observed as it decreases from 0.2 to about 0.1. Further enlargement of the training data set does not alter much the testing loss, which remains at the level of 0.1. Hence, 20 groups of the training DNS data are adopted in the current work.
Figure 5: The evolutions of the testing loss with different amount of training data.
The training and testing losses for $Re_{\tau}=395$ and $590$ are shown in Figs. 6a and 6b, respectively. It can be seen that as $Re_{\tau}$ increases, the loss goes up, indicating the increased difficulty in learning high Reynolds number turbulent flow. Meanwhile, to avoid over-fitting of the model, we have applied early stopping which is widely used in gradient descent learning [105, 106]. In both cases, the model parameters are extracted after 30 epochs before the performance on the testing set starts to degrade.
@@
In Figs. 7, 8 and 9, the mean streamwise velocity and root-mean-squared (rms) fluctuating velocities predicted by the IUFNO model are displayed for $Re_{\tau}\approx 180$, $395$ and $590$, respectively. Also shown in the figures are the corresponding predictions by the DSM and WALE models. As can be seen, the IUFNO model performs reasonably better at all three Reynolds numbers compared to the DSM and WALE models. Nevertheless, the prediction of the mean velocity by IUFNO at $Re_{\tau}\approx 590$ is not as good as that at $Re_{\tau}\approx 180$ and $395$. This also reflects the increasing difficulty in predicting turbulent flows at high Reynolds number. For the traditional LES models, we observe that the WALE model performs tangibly better than the DSM. Meanwhile, the WALE model diverges at $Re_{\tau}\approx 590$, thus the results of WALE are not presented for $Re_{\tau}\approx 590$. Here, it should be emphasized that the current LES grids are coarser than the general requirements for wall-resolved LES [5, 102]. In the expense of additional computational cost (i.e. using sufficiently fine grids or carefully configured wall model), the traditional LES models can also achieve reasonably accurate LES predictions, which is not the concern of the current work.
@@
The probability density functions (PDFs) of the three fluctuating velocity components are shown in Fig. 10 for different friction Reynolds numbers. The PDFs of the velocity fluctuations are all symmetric, and the ranges of the PDFs are in consistency with the corresponding rms values in Figs. 7, 8 and 9. It can be seen that the IUFNO model gives reasonably good predictions of the PDFs at all three Reynolds numbers. Notably, the advantage of the IUFNO model is more pronounced for the streamwise components of the fluctuating velocity compared to the normal and spanwise components whose predictions by the WALE model are also satisfying. The DSM gives the worst predictions among all models.
Figure 10: The probability density functions (PDFs) of the velocity fluctuation components at different Reynolds numbers: (a) $u^{\prime},Re_{\tau}\approx 180$; (b) $v^{\prime},Re_{\tau}\approx 180$; (c) $w^{\prime},Re_{\tau}\approx 180$; (d) $u^{\prime},Re_{\tau}\approx 395$; (e) $v^{\prime},Re_{\tau}\approx 395$; (f) $w^{\prime},Re_{\tau}\approx 395$; (g) $u^{\prime},Re_{\tau}\approx 590$; (h) $v^{\prime},Re_{\tau}\approx 590$; (i) $w^{\prime},Re_{\tau}\approx 590$.
The predicted shear Reynolds stresses by different LES models are shown in Fig. 11. The maximum shear Reynolds stresses are located near the upper and lower walls where both the mean shear effects and the velocity fluctuations are strong. The distribution between the two peaks is approximately linear, which is consistent with the literature [107]. Both the IUFNO and WALE models can predict the Reynolds stress well at $Re_{\tau}\approx 180$ and $395$ while apparent discrepancy can be observed for the DSM. At $Re_{\tau}\approx 590$, the IUFNO model still performs better than the DSM, but it has some wiggles in the linear region.
@@
As shown in Fig. 12, at all three Reynolds numbers, the joint PDFs close to the wall are much wider in the second and fourth quadrants, indicating that shear events (ejection and sweep) are more dominant near the wall rather than the center of the channel. It can be seen that the IUFNO results agree well with the fDNS benchmark. Meanwhile, the predictions by the WALE and DSM models are also satisfying.
@@
To examine the energy distribution at different scales, we calculate the kinetic energy spectrum in the LES. The streamwise spectra are shown in Figs. 13a, 13b and 13c for $Re_{\tau}\approx 180$, $395$ and $590$, respectively. At $Re_{\tau}\approx 180$, the predicted spectrum by the IUFNO model agrees well with the fDNS result. Both the DSM and WALE models overestimate the energy spectrum while the WALE model outperforms the DSM. This is also consistent with the results for the rms fluctuating velocity magnitude in Fig. 7. At $Re_{\tau}\approx 395$, the spectrum predicted by IUFNO slightly deviates from the fDNS result, but it still outperforms the DSM and WALE models. For the spectrum at $Re_{\tau}\approx 590$, the IUFNO model still outperforms the DSM, however, some nonphysical discontinuities in the IUFNO spectrum can be observed. The cause of such phenomenon is yet unclear.
The spanwise kinetic energy spectra are shown in Figs. 13d, 13e and 13f for $Re_{\tau}\approx 180$, $395$ and $590$, respectively. Unlike the streamwise spectrum, the maximum energy in the spanwise spectrum is not contained in the first wavenumber, presumably because there are no dominant largest-scale motions in the spanwise direction. Overall, the IUFNO model has a closer agreement with the fDNS result compared to the WALE and DSM models.
@@
where $\overline{\Omega}_{ij}=\frac{1}{2}(\partial{\overline{u}_{i}}/\partial{x_{j}}-\partial{\overline{u}_{j}}/\partial{x_{i}})$ is the filtered rotation rate. The instantaneous iso-surfaces of $Q$ for $Re_{\tau}\approx 180$, $395$ and $590$ are displayed in Figs. 14, 15 and 16, respectively. The iso-surfaces are colored by the streamwise velocities. As the figures depict, the IUFNO models have overall closer agreements with the fDNS results compared to the traditional LES models. Hence, the ability of the IUFNO model to predict the vortex structures is confirmed.
@@
Finally, we compare the computation cost in the LES for different models in Table 3. The values in the table are the time consumption (in seconds) for equivalent 10000 DNS time steps. The IUFNO models are tested on the NVIDIA A100 GPU, where the CPU type is AMD EPYC 7763 @2.45GHz. The DSM and WALE simulations are performed on the CPU, which is Intel Xeon Gold 6148 @2.40 GHz. As can be seen in Table 3, among the FNO-based models, the original FNO has the lowest computational cost. IUFNO has the largest computational cost as it combines the structures of both the IFNO and UFNO. Nevertheless, the computational efficiency of the IUFNO model is still considerably higher than those of the traditional LES models (i.e. DSM and WALE), considering that the runtimes of the traditional models are not multiplied by the number of adopted computation cores (shown in the parentheses).
$Re_{\tau}$ DSM WALE FNO IFNO UFNO IUFNO
180 97.4s ($\times$16 cores) 52.1s ($\times$16 cores) 1.3s 2.9s 1.6s 5.0s
395 287.2s ($\times$32 cores) 146.3s ($\times$32 cores) 4.5s 9.1s 5.3s 12.1s
590 315.3s ($\times$64 cores) NA 5.9s 11.6s 6.7s 14.9s
@@
Apparently, the current work shows the potential of IUFNO in the long-term prediction of turbulent channel flow. Nevertheless, it is worth noting that there are still many challenges for machine-learning-based flow predictions, such as high Reynolds number turbulence, turbulent flows with complex geometries, irregular and non-Cartesian grids, etc. On the other hand, it is noticeable that the neural operator-based methods are also quickly evolving and improving these days [55, 65, 79, 89]. For example, incorporating physical constraints to FNO can lower the required amount of the training data [89]. The recently proposed geometry-informed neural operator (GINO) can handle complex geometries by combining the FNO with graph neural operator (GNO) which can transform irregular grids into uniform ones [79]. Meanwhile, the transformer-based frameworks and its combination with FNO have been developed recently, which are also potentially suitable for complex flows [70, 88]. These are several new aspects that deserve future investigations for complex turbulent flows at high Reynolds numbers.